# machine learning UWB ultra-wideband

### v1

@@ Abstract
Index Terms:
Foundation models, LLMs, wireless networks
@@ I Introduction
In the ever-evolving landscape of wireless communications and networking research, including the advancements of the upcoming 6G standard, artificial intelligence (AI) has become the focal point for tackling issues that cannot be solved by traditional non-machine learning approaches [1]. Such approaches are based on traditional optimization theory techniques and only work if suitable mathematical models are available. With increasing user traffic and density, formulating such models is difficult [2]. Deep learning is suitable in such situations because it can find effective functions from a dataset, while aiming to be scalable and generalizable. It powers a wide range of applications such as recognizing wireless technologies, wireless optimization, classifying non-line-of-sight (NLOS) conditions, estimating environmental parameters and analyzing human behavior in sports and healthcare applications. [3, 4, 5]. Recently, significant efforts have been made by 3rd Generation Partnership Project (3GPP) to standardize machine learning (ML) in wireless networks. 3GPP has integrated ML into the 5G core architecture in releases 15 and 16, while ’Open RAN’ embraces a ML-native approach [6].
@@
    In the first phase of the development of wireless AI (see Figure 1), many researchers individually proposed ML models for different wireless communication tasks such as recognition of wireless technologies, management of radio resources, classification of (N)LOS, etc., which require domain expertise and expensive ML model training costs [7]. Although these models use popular architectures such as deep, convolutional, and long short-term memory neural networks, the idea of sharing knowledge (weights) across various wireless downstream tasks is vastly underexplored.
@@
    A second phase of advances in AI has been led by domains that have a lot of data available, such as neurolinguistic programming (NLP) and the computer vision domain. In such domains, neural networks have been thoroughly investigated and successfully demonstrated over the last decade. During this period, transfer learning has been proposed to leverage and synthesize the knowledge distilled within a task and use valuable experience accumulated in the past to facilitate the learning of new problems [8]. This requires less labeled data, ML can adapt quickly to new environments and can be more robust. However, this approach focuses on variations within single tasks, including those in the wireless domain.
@@
    Very recently, in the third phase, key AI domains have undergone a paradigm shift with the introduction of foundation models. A foundation model is a ”paradigm for building AI systems” in which a model trained on a large amount of unlabeled data can be adapted to many different downstream tasks, even with different types of data [9]. For example, a foundation model for image generation can cope with a wide variety of styles and shapes. Upon showing just a few new samples of a new style, the foundation model is adapted towards a new downstream task (e.g., generation of images using the new style) without significant effort required. When we consider these rapid advances in AI, it becomes evident that current task-based ML approaches in wireless networks are no longer the way forward.
@@
In the next section, we explore the advances in foundation models and discuss the need for WPFMs to streamline AI development in various tasks in wireless communication. Next, Section III presents the challenges in the wireless domain of building a WPFM. Section IV outlines wireless applications, used to demonstrate the foundation model in Section V, where we present a general WPFM framework and implementation strategies. Next, Section VI discusses the next research steps for WPFMs and enabling human-interactive wireless networks. Finally, conclusions are drawn in Section VII.
@@ II Advances in foundation models
Foundation models have gained a lot of interest in the field of natural language processing and computer vision with the introduction of models such as GPT3/4, ChatGPT, BERT, DALL-E, etc. [10]. These models have enabled significant progress and achieved state of the art (SOTA) accuracy in their fields. With the abundant availability of large datasets, foundation models can be seamlessly applied to a diverse set of downstream tasks. Innovations to such foundation models include unsupervised pre-training techniques that enables models to grasp complex patterns and relationships within massive datasets. Additionally, fine-tuning strategies have improved in recent years, e.g. few- and zero-shot learning for new applications.
To accommodate diverse data types, multimodality has become a prominent feature in machine learning and is frequently adapted by foundation models [11]. Models such as ’CLIP’, designed for vision tasks, seamlessly merge text and image encodings, facilitating zero-shot image identification [12]. Additionally, text-to-image models like DALL-E demonstrate a profound synergy between language and vision models, incorporating encoded text prompts and image generation capabilities [13]. The autoregressive-trained transformer can simultaneously encode text and images, showcasing robust performance even in zero-shot evaluation scenarios. This integration of modalities not only enhances versatility, but also underscores the adaptability of (transformer-based) neural network architectures to handle varied input types effectively.
@@
One domain facing this challenge, i.e. robotics, has started to embrace the collection large dataset to achieve a general-purpose (foundation) model [14]. Instead of relying on single-task AI, which only works well on a single robot involved in a single experiment, researchers have collaborated on achieving a general AI robot that has learned from a pool of many robots. Surprisingly, the authors found that multirobot data could be used with similar methods used in LLMs, provided that they follow the recipe of using large neural networks with large datasets. The dataset contains five hundred different skills and interactions with thousands of different objects, and allows training of a foundation model without requiring any special features for cross-embodiment. Remarkably, the global (foundation) model outperformed the robots trained in individual tasks. Finally, the authors have successfully integrated large vision and language models, which are finetuned to describe actions based on visual data from the robot. This allows prompt-based robotic actions, even with unseen objects.
@@
In addressing the challenges posed by the advent of 6G networks, encompassing Non-Terrestrial Networks (NTN), Terrestrial Networks (TN), mmWave/THz network, underwater networks, and advancements in Wireless Local Area Networks (WLANs) [15], could a unified WPFM serve as a viable solution to streamline and reduce redundancy in AI development across diverse wireless networks? Much like in the domain of robotics, similar challenges exist, where large datasets sourced from the Internet are scarce or too simplistic to represent complex wireless conditions. However, the availability of data is quickly catching up. Popular journals are soliciting papers specific to datasets [16], as depicted in Figure 1.
We believe it is crucial to start the development of a foundation model which focuses on the wireless physical layer. Unlike higher layers in wireless networks, the realization of this WPFM requires dealing with complex and dynamic data in the form of time series. The adoption of ML innovations can be significantly accelerated with a model that can understand and represent these time series. Trained foundation models using architectures such as transformer networks will eliminate the need for exhaustive research, domain expertise, expensive data collection, as it enables improved representation learning across multiple downstream tasks.
@@
Next, it will be possible to integrate a WPFM with foundation models intended for higher layers of the network stack. In these layers, pre-trained LLMs, such as ChatGPT, BERT, and Llama, have been very recently investigated and can already perform general tasks quite well [9] [17] [18]. Examples of such tasks include answering healthcare-related questions, classifying 3GPP working group based on technical specifications, answering telecom-related questions, etc. However, these models lack an understanding of physical wireless signals. With a WPFM, we can integrate existing LLMs to open up a semantic understanding of the physical wireless environment. System operators will no longer solely rely on traditional rule-based configurations programmed by engineers for individual tasks (e.g. defined channel selection and power allocation parameters for each application); they will be able to simply instruct (prompt) the model to configure networks or build applications using in-context learning (giving input and output examples) and utilize automatic optimization (without explicit programming) using physical-layer sensing WPFMs, as illustrated in Figure 1.
@@ III Challenges in wireless physical-layer foundation models
Recent scientific work has provided preliminary building blocks for WPFMs, such as transfer learning solutions and per-task data representations. However, since there is no general methodology to integrate these building blocks into an overall foundation model, significant research remains to be done. In addition, the usage of LLMs has been a recent research topic and even standardization within wireless communications, however, no physical-layer results have been presented and fundamental research challenges have yet to be identified [17, 18, 19, 20]. This paper addresses the following fundamental research challenges:
@@ III-1 Challenge 1: Novel methods are required for wireless pre-training tasks
Modern ML architectures are capable of handling complex pattern learning and scaling of many parameters. The transformer neural network [21], for example, excels in efficiently learning sequential data with parallel training. In other domains, foundation models exist which can perform general tasks, such as CLIP, DALL-E and LLMs such as GPT3, GPT4 and BERT, but also have been applied recently in time series [22, 23]. However, in wireless networks, the transformer architecture is missing its potential by being trained to perform individual tasks. Instead, training large foundation transformer models requires effective un- and self-supervised pre-training tasks, e.g., in LLMs next-word predictions, masking, positional predictions, denoising, etc., are used [9]. This is key in enabling a shared foundation model for multiple downstream tasks (Figure 1). However, with limited data availability in the past and challenges in supporting heterogeneous data types, there is a gap in the wireless domain for effective pre-training tasks.
@@
Creating a foundation model for wireless time series, such as in-phase and quadrature (IQ) samples and channel impulse responses, diverges significantly from models in the NLP (i.e. semantic understanding) and vision domains (i.e. 2D / 3D patterns) due to the heterogeneous temporal nature of the data. Unlike text or images, time-series data unfold over time, which presents challenges related to capturing embedding temporal dependencies, modeling variations, and addressing issues like fading channels and dynamic network conditions. Another contrast with SOTA can be found in forecasting time series. Here, the focus is on predicting future trends based on historical data and seasonal trends, often seen in financial markets or climate studies, emphasizing precision and accuracy in predicting future outcomes. Analyzing wireless networks also proves to be much more difficult when dealing with heterogeneous wireless time series which are non-intuitive, in contrast to data in the aforementioned domains. Wireless SOTA embedding is limited to reducing the dimensionality (compression) or mapping (in)complete data in a homogeneous space [8]. Developing a WPFM demands a deep understanding of time-sensitive data dynamics heterogeneous, i.e. supporting samples from use cases with different lengths, sampling rates and data types (IQ, received signal strength indicator (RSSI), CIR, metadata etc.), setting it apart as a compelling and distinctive domain within the broader AI landscape.
@@
The semantic description of wireless networks and their environments has been a recent topic for standalone (wireless) machine learning models [12, 24]. Combining textual descriptions with the first two challenges in a foundation model allows AI to be used in many downstream applications and opens up new possibilities. For example, by combining the embedded information from WPFMs with such semantic descriptions, we allow LLMs to autonomously understand wireless data from the physical layer, enable AI chain-of-thoughts and prompt-based applications and optimization. SOTA lacks the ability for wireless optimization (LLM + WPFM) and is currently limited to LLM-based configurations using standardization and specification documents [25].
@@ V A wireless physical-layer foundation model strategic framework
To address the research challenges, we present a strategic framework of a WPFM. This foundation model, illustrated in Figure 2 has a simple, but transformative goal: it can understand many types of physical wireless signals (time series) and metadata, while enabling network configurations and optimization and allowing user interaction. This will enable easy transfer to new downstream tasks (e.g., CIR understanding to detect different types of objects using UWB radar, IQ pattern encoding to recognize unseen wireless technologies in a spectrum manager and configuring a 6G network depending on the environmental conditions such as interfering signals, noise, etc.).
@@ V-A Embedding and tokenization of physical-layer time-series
In the wireless domain, technologies and applications generally consist of heterogeneous time series. This means that the length of the input data, its data type and dimensionality can vary across applications. To ensure a consistent input for foundation models, we propose using embedding or tokenizing of wireless time series data, which has several implementation options. One possibility is to use Byte-Pair Encoding (BPE) to merge frequently occurring signal or pulse IQ sample pairs, until a fixed model input size is obtained. Another possibility is to represent known pulses and short-term patterns in time series as tokens. Alternatively, Time Variable based Tokenization (TVT) can be used to tokenize time series where the token uniformity inductive bias would act on the variable dimension [26]. In addition to encoding, it is essential to define the vocabulary size (number of unique tokens) to support variable input lengths, without making the model unnecessary complex. By applying tokenization, we can provide the foundation model with a uniform sample distribution, irrespective of the sampling rate (reaching e.g., <<< 10% embedding difference) and is robust to input noise (e.g., SNR >>> 10 dB). Furthermore, it should support samples at different bandwidths or sample rates and allow generative AI of samples close (in embedded space) to a wireless time series for data augmentation or simulation purposes. The result of a robust tokenizer can serve as input for a foundation model based on, for example, transformers.
@@ V-C Semantic physical-layer wireless representation learning
Use case applications Metadata such as the environment dimensions (walls, ceiling height, etc.), type of environment (outdoor, office indoor, industrial environment, etc.), number of obstacles, persons, etc. can be included, as it impacts the efficiency and performance of wireless applications. Such descriptions can be augmented by LLMs and allow foundation models to generatively provide context to new, unknown environments. Additionally, the generative capabilities can help with diagnostics or simulations of UWB and wireless deployments.

### v2

@@ I Introduction
In the ever-evolving landscape of wireless communications and networking research, including the advancements of the upcoming 6G standard, artificial intelligence (AI) has become the focal point for tackling issues that cannot be solved by traditional non-machine learning approaches [9446676]. Such approaches are based on traditional optimization theory techniques and only work if suitable mathematical models are available. With increasing user traffic and density, formulating such models is difficult [zappone2019wireless]. Deep learning is suitable in such situations because it can find effective functions from a dataset, while aiming to be scalable and generalizable. It powers a wide range of applications such as recognizing wireless technologies, wireless optimization, classifying non-line-of-sight (NLOS) conditions, estimating environmental parameters and analyzing human behavior in sports and healthcare applications. [fontaine2019towards, cheng2021deep, fontaine2023transfer]. Recently, significant efforts have been made by 3rd Generation Partnership Project (3GPP) to standardize machine learning (ML) in wireless networks. 3GPP has integrated ML into the 5G core architecture in releases 15 and 16, while ‘Open RAN’ embraces a ML-native approach [9768336].
@@
    In the first phase of the development of wireless AI (see Figure 1), many researchers individually proposed ML models for different wireless communication tasks such as recognition of wireless technologies, management of radio resources, classification of (N)LOS, etc., which require domain expertise and expensive ML model training costs [kulin2021survey]. Although these models use popular architectures such as deep, convolutional, and long short-term memory neural networks, the idea of sharing knowledge (weights) across various wireless downstream tasks is vastly underexplored.
@@
    A second phase of advances in AI has been led by domains that have a lot of data available, such as neurolinguistic programming (NLP) and the computer vision domain. In such domains, neural networks have been thoroughly investigated and successfully demonstrated over the last decade. During this period, transfer learning has been proposed to leverage and synthesize the knowledge distilled within a task and use valuable experience accumulated in the past to facilitate the learning of new problems [nguyen2021transfer]. This requires less labeled data, ML can adapt quickly to new environments and can be more robust. However, this approach focuses on variations within single tasks, including those in the wireless domain.
@@
    Very recently, in the third phase, key AI domains have undergone a paradigm shift with the introduction of foundation models. A foundation model is a “paradigm for building AI systems” in which a model trained on a large amount of unlabeled data can be adapted to many different downstream tasks, even with different types of data [zhou2023comprehensive]. For example, a foundation model for image generation can cope with a wide variety of styles and shapes. Upon showing just a few new samples of a new style, the foundation model is adapted towards a new downstream task (e.g., generation of images using the new style) without significant effort required. When we consider these rapid advances in AI, it becomes evident that current task-based ML approaches in wireless networks are no longer the way forward.
@@
In the next section, we explore the advances in foundation models and discuss the need for WPFMs to streamline AI development in various tasks in wireless communication and compare its difference with LLMs. Next, Section III presents the challenges in the wireless domain of building a WPFM. Section IV outlines wireless applications, used to demonstrate the foundation model in Section V, where we present a general WPFM framework and implementation strategies. Next, Section VI discusses the next research steps for WPFMs and enabling human-interactive wireless networks. Finally, conclusions are drawn in Section VII.
@@ II Advances in foundation models
II-A The rise of foundation models
Foundation models have gained a lot of interest in the field of natural language processing and computer vision with the introduction of models such as GPT3/4, ChatGPT, BERT, DALL-E, etc. [brown2020language]. These models have enabled significant progress and achieved state of the art (SOTA) accuracy in their fields. With the abundant availability of large datasets, foundation models can be seamlessly applied to a diverse set of downstream tasks. Innovations to such foundation models include unsupervised pre-training techniques that enables models to grasp complex patterns and relationships within massive datasets. Additionally, fine-tuning strategies have improved in recent years, e.g. few- and zero-shot learning for new applications.
To accommodate diverse data types, multimodality has become a prominent feature in machine learning and is frequently adapted by foundation models [liu2022zero]. Models such as ‘CLIP’, designed for vision tasks, seamlessly merge text and image encodings, facilitating zero-shot image identification [radford2021learning]. Additionally, text-to-image models like DALL-E demonstrate a profound synergy between language and vision models, incorporating encoded text prompts and image generation capabilities [ramesh2021zero]. The autoregressive-trained transformer can simultaneously encode text and images, showcasing robust performance even in zero-shot evaluation scenarios. This integration of modalities not only enhances versatility, but also underscores the adaptability of (transformer-based) neural network architectures to handle varied input types effectively.
@@ II-A The rise of foundation models
One domain facing this challenge, i.e. robotics, has started to embrace the collection large dataset to achieve a general-purpose (foundation) model [padalkar2023open]. Instead of relying on single-task AI, which only works well on a single robot involved in a single experiment, researchers have collaborated on achieving a general AI robot that has learned from a pool of many robots. Surprisingly, the authors found that multirobot data could be used with similar methods used in LLMs, provided that they follow the recipe of using large neural networks with large datasets. The dataset contains five hundred different skills and interactions with thousands of different objects, and allows training of a foundation model without requiring any special features for cross-embodiment. Remarkably, the global (foundation) model outperformed the robots trained in individual tasks. Finally, the authors have successfully integrated large vision and language models, which are finetuned to describe actions based on visual data from the robot. This allows prompt-based robotic actions, even with unseen objects.
@@
In addressing the challenges posed by the advent of 6G networks, encompassing Non-Terrestrial Networks (NTN), Terrestrial Networks (TN), mmWave/THz network, underwater networks, and advancements in Wireless Local Area Networks (WLANs) [10.1145/3571072], could a unified WPFM serve as a viable solution to streamline and reduce redundancy in AI development across diverse wireless networks? Much like in the domain of robotics, similar challenges exist, where large datasets sourced from the Internet are scarce or too simplistic to represent complex wireless conditions. However, the availability of data is quickly catching up. Popular journals are seeking papers specific to datasets [fischione2023data], as depicted in Figure 1.
II-B Strategic differences between LLMs and WPFMs
@@
LLMs have already been popularized in wireless networks with examples of tasks that include answering healthcare-related questions, classifying 3GPP working group based on technical specifications, answering telecom-related questions, network configurations, etc. [bariah2023large] [tong2023ten]. However, since these models lack an understanding of physical wireless signals, we believe the development of a WPFM is crucial and opens up a semantic understanding of the physical wireless environment. Unlike text-based configurations in the higher layers of wireless networks, the realization of a WPFM requires dealing with complex and dynamic data in the form of time series. The adoption of ML innovations can be significantly accelerated with a model that can understand and represent these time series. Trained foundation models using architectures such as transformer networks will eliminate the need for exhaustive research, domain expertise, and expensive data collection, as it enables improved representation learning across multiple downstream tasks. Next, it will be possible to integrate a WPFM with LLMs which focus on higher layers of the network stack. In these layers, pre-trained LLMs, such as ChatGPT, BERT, and Llama, have been very recently investigated and can already perform general tasks quite well [zhou2023comprehensive] [lin2023pushing] [maatouk2023large]. System operators will no longer solely rely on traditional rule-based configurations programmed by engineers for individual tasks (e.g. defined channel selection and power allocation parameters for each application); they will be able to simply instruct (prompt) LLMs to configure networks or build applications using in-context learning (giving input and output examples) and utilize automatic optimization (without explicit programming) with physical-layer sensing WPFMs, as illustrated in Figure 1.
@@ III Challenges in wireless physical-layer foundation models
Recent scientific work has provided preliminary building blocks for WPFMs, such as transfer learning solutions and per-task data representations. However, since there is no general methodology to integrate these building blocks into an overall foundation model, significant research remains to be done. In addition, the usage of LLMs has been a recent research topic and even standardization within wireless communications, however, no physical-layer results have been presented and fundamental research challenges have yet to be identified [lin2023pushing, maatouk2023large, bariah2023large, tong2023ten]. This paper addresses the following fundamental research challenges:
@@ III-1 Challenge 1: Novel methods are required for wireless pre-training tasks
Modern ML architectures are capable of handling complex pattern learning and scaling of many parameters. The transformer neural network [vaswani2017attention], for example, excels in efficiently learning sequential data with parallel training. In other domains, foundation models exist which can perform general tasks, such as CLIP, DALL-E and LLMs such as GPT3, GPT4 and BERT, but also have been applied recently in time series [wang2022transformer, wang2022wir]. However, in wireless networks, the transformer architecture is missing its potential by being trained to perform individual tasks. Instead, training large foundation transformer models requires effective un- and self-supervised pre-training tasks, e.g., in LLMs next-word predictions, masking, positional predictions, denoising, etc., are used [zhou2023comprehensive]. This is key in enabling a shared foundation model for multiple downstream tasks (Figure 1). However, with limited data availability in the past and challenges in supporting heterogeneous data types, there is a gap in the wireless domain for effective pre-training tasks.
@@
Creating a foundation model for wireless time series, such as in-phase and quadrature (IQ) samples and channel impulse responses, diverges significantly from models in the NLP (i.e. semantic understanding) and vision domains (i.e. 2D / 3D patterns) due to the heterogeneous temporal nature of the data. Unlike text or images, time-series data unfold over time, which presents challenges related to capturing embedding temporal dependencies, modeling variations, and addressing issues like fading channels and dynamic network conditions. Another contrast with SOTA can be found in forecasting time series. Here, the focus is on predicting future trends based on historical data and seasonal trends, often seen in financial markets or climate studies, emphasizing precision and accuracy in predicting future outcomes. Analyzing wireless networks also proves to be much more difficult when dealing with heterogeneous wireless time series which are non-intuitive, in contrast to data in the aforementioned domains. Wireless SOTA embedding is limited to reducing the dimensionality (compression) or mapping (in)complete data in a homogeneous space [nguyen2021transfer]. Developing a WPFM demands a deep understanding of time-sensitive data dynamics heterogeneous, i.e. supporting samples from use cases with different lengths, sampling rates and data types (IQ, received signal strength indicator (RSSI), CIR, metadata etc.), setting it apart as a compelling and distinctive domain within the broader AI landscape.
@@
The semantic description of wireless networks and their environments has been a recent topic for standalone (wireless) machine learning models [radford2021learning, li2019visualbert]. Combining textual descriptions with the first two challenges in a foundation model allows AI to be used in many downstream applications and opens up new possibilities. For example, by combining the embedded information from WPFMs with such semantic descriptions, we allow LLMs to autonomously understand wireless data from the physical layer, enable AI chain-of-thoughts and prompt-based applications and optimization. SOTA lacks the ability for wireless optimization (LLM + WPFM) and is limited to LLM-based configurations using standardization and specification documents [kotaru2023adapting].
@@ V A wireless physical-layer foundation model strategic framework
To address the research challenges, we present a strategic framework of a WPFM. This foundation model, illustrated in Figure 2 has a simple, but transformative goal: it can understand many types of physical wireless signals (time series) and metadata, while enabling network configurations and optimization and allowing user interaction. Using this framework and large openly available datasets, researchers can enable the usage of WPFM by companies and/or new researchers in the field to easily adapt WPFMs to new downstream tasks (e.g., CIR understanding to detect different types of objects using UWB radar, IQ pattern encoding to recognize unseen wireless technologies in a spectrum manager and configuring a 6G network depending on the environmental conditions such as interference signals, noise, etc.).
@@ V-A Embedding and tokenization of physical-layer time-series
In the wireless domain, technologies and applications generally consist of heterogeneous time series. This means that the length of the input data, its data type and dimensionality can vary across applications. To ensure consistent input for foundation models, we propose using embedding or tokenizing of wireless time series data, which has several implementation options. One strategy is to use Byte-Pair Encoding (BPE) to merge frequently occurring signal or pulse IQ sample pairs, until a fixed input size of the model is obtained. Another strategy is to represent known pulses and short-term patterns in time series as tokens. Alternatively, Time Variable based Tokenization (TVT) can be used to tokenize time series where the token uniformity inductive bias would act on the variable dimension [zhou2022k]. In addition to encoding, it is essential to define the vocabulary size (number of unique tokens) to support variable input lengths while avoiding unnecessary complexity. For instance, by applying tokenization techniques such as those utilized in models like BERT or GPT, we can ensure a uniform sample distribution regardless of the sampling rate. This results in a similar embedding space, e.g. << 10% difference, even when comparing signals sampled at different rates (e.g. 10 or 20 MHz) or when subjected to input noise with e.g. a signal-to-noise ratio (SNR) >> 10 dB. Such tokenization strategies are particularly valuable in scenarios involving signal processing at different bandwidths or sample rates and support data augmentation or simulation tasks. For example, they enable the generation of synthetic samples closely resembling wireless time series data, facilitating realistic simulations for various applications. The result is a robust tokenizer that can be integrated in well-defined foundation models such as transformers.
@@ V-C Semantic physical-layer wireless representation learning
Use case applications Metadata such as the dimensions of the environment (walls, ceiling height, etc.), type of environment (outdoor, office indoor, industrial environment, etc.), number of obstacles, persons, etc. can be included, as it impacts the efficiency and performance of wireless applications. Such descriptions can be augmented by LLMs and allow foundation models to generatively provide context to new, unknown environments. Additionally, generative capabilities can help with diagnostics or simulations of UWB and wireless deployments.